Find the exact value of sin theta if, theta = tan inverse (2/3)
Answer by lwsshak3(11628)

Find the exact value of sin theta if, theta = tan inverse (2/3)
***
tan theta=2/3
hypotenuse 0f working right triangle=√(2^2+3^2)=√(4+9)=√13
sin theta=2/√13=2√13/13
check:(w/calculator)
tan x=2/3
x≈33.69˚
sinx=sin33.69˚≈0.5547..
exact value as calculated=2√13/13≈0.5547